# Interaction-induced phase transitions and critical phases in nonreciprocal non-Hermitian quasicrystals

### v1

@@ Abstract
Non-Hermitian phenomena, such as exceptional points, non-Hermitian skin effects, and topologically nontrivial phases, have attracted continued attention. In this work, we reveal how interactions and nonreciprocal hopping could collectively influence the behavior of two interacting bosons on quasiperiodic lattices. Focusing on the Bose-Hubbard model with Aubry-André-Harper quasiperiodic modulations and hopping asymmetry, we discover that the interaction could enlarge the localization transition point of the noninteracting system into a critical phase, in which localized doublons formed by bosonic pairs can coexist with delocalized states. Under the open boundary condition, the bosonic doublons could further show non-Hermitian skin effects, realizing doublon condensation at the edges, and their direction of skin-localization can be flexibly tuned by the hopping parameters. A framework is developed to characterize the spectral, localization, and topological transitions accompanying these phenomena. Our work advances the understanding of localization and topological phases in non-Hermitian systems, particularly in relation to multiparticle interactions.
@@ I Introduction
Non-Hermitian systems, usually characterized by non-Hermitian Hamiltonians constitute ideal setups to explore the interplay among localization, environmental effects and interactions. In recent years, non-Hermitian systems have attracted increasing attention due to their distinctive physics, such as the parity-time-reversal (PT) transitions [20, 21, 22], exceptional points [23, 24, 25, 26, 27, 28], non-Hermitian skin effects (NHSEs) [29, 30, 31, 32, 33] and enriched symmetry classifications of topological phases [34, 35, 36, 37, 38, 39, 40]. Meanwhile, the interplay between disorder and non-Hermitian effects may lead to unique types of spectral, localization, topological and entanglement transitions [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74], yielding universality classes of localization beyond closed-system contexts [75, 76, 77, 78]. With additional elements like lattice dimerization, long-range hopping, non-Abelian potential and time-period driving, critical phases and mobility edges could also emerge in disordered non-Hermitian systems even in one spatial dimension [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96], enriching further the non-Hermitian localization phenomenon.
Interaction could greatly complicates the AL in non-Hermitian systems due to its active competition with other localization mechanisms, including random or correlated disorder and NHSEs. Great efforts have been made towards the understanding of “many-body localization” in non-Hermitian settings, with a focus on multiple interacting particles in lattices of relatively small sizes [97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. Nevertheless, considering a few interacting particles in relatively large lattices may offer necessary insights regarding how interactions could drive different phase transitions and criticality in the thermodynamic limit (with system size $L\rightarrow\infty$) [108, 109, 110, 111, 112, 113, 114, 115, 116]. Motivated by these considerations, we explore in this work the interaction-driven phase transitions and critical phases in non-Hermitian quasicrystals (NHQCs). Focusing on two interacting bosons in one-dimensional (1D) Aubry-André-Harper (AAH) lattices with quasiperiodic onsite potentials and nonreciprocal hoppings, we reveal that the interactions could not only transform the AL critical points of noninteracting systems into critical phases, but also push the boundaries of fully localized phases towards stronger asymmetric hopping regions. Within a critical phase, extended two-boson states can coexist with bounded boson pairs forming doublons, which could also subject to NHSEs under the open boundary condition (OBC). The borders between extended, critical and localized phases are further found to carry topologically nontrivial signatures.

### v2

@@ Abstract
Non-Hermitian phenomena, such as exceptional points, non-Hermitian skin effects, and topologically nontrivial phases have attracted continued attention. In this work, we reveal how interactions and nonreciprocal hopping could collectively influence the behavior of two interacting bosons on quasiperiodic lattices. Focusing on the Bose-Hubbard model with Aubry-André-Harper quasiperiodic modulations and hopping asymmetry, we discover that interactions could enlarge the localization transition point of the noninteracting system into a critical phase, in which localized doublons formed by bosonic pairs can coexist with delocalized states. Under the open boundary condition, the bosonic doublons could further show non-Hermitian skin effects, realizing doublon condensation at the edges, and their direction of skin-localization can be flexibly tuned by the hopping parameters. A framework is developed to characterize the spectral, localization, and topological transitions accompanying these phenomena. Our work advances the understanding of localization and topological phases in non-Hermitian systems, particularly in relation to multiparticle interactions.
@@ I Introduction
Non-Hermitian systems, usually characterized by non-Hermitian Hamiltonians constitute ideal setups to explore the interplay among localization, environmental effects and interactions. In recent years, non-Hermitian systems have attracted increasing attention due to their distinctive physics, such as the parity-time-reversal (PT) transitions [20, 21, 22], exceptional points [23, 24, 25, 26, 27, 28], non-Hermitian skin effects (NHSEs) [29, 30, 31, 32, 33, 34] and enriched symmetry classifications of topological phases [35, 36, 37, 38, 39, 40, 41]. Meanwhile, the interplay between disorder and non-Hermitian effects may lead to unique types of spectral, localization, topological and entanglement transitions [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75], yielding universality classes of localization beyond closed-system contexts [76, 77, 78, 79]. With additional elements like lattice dimerization, long-range hopping, non-Abelian potential and time-period driving, critical phases and mobility edges could also emerge in disordered non-Hermitian systems even in one spatial dimension [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97], enriching further the non-Hermitian localization phenomenon.
Interaction could greatly complicates the AL in non-Hermitian systems due to its active competition with other localization mechanisms, including random or correlated disorder and NHSEs. Great efforts have been made towards the understanding of “many-body localization” in non-Hermitian settings, with a focus on multiple interacting particles in lattices of relatively small sizes [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. Nevertheless, considering a few interacting particles in relatively large lattices may offer necessary insights regarding how interactions could drive different phase transitions and criticality in the thermodynamic limit (with system size $L\rightarrow\infty$) [109, 110, 111, 112, 113, 114, 115, 116, 117, 118]. Motivated by these considerations, we explore in this work the interaction-driven phase transitions and critical phases in non-Hermitian quasicrystals (NHQCs). Focusing on two interacting bosons in one-dimensional (1D) Aubry-André-Harper (AAH) lattices with quasiperiodic onsite potentials and nonreciprocal hoppings, we reveal that the interactions could not only transform the AL critical points of noninteracting systems into critical phases, but also push the boundaries of fully localized phases towards stronger asymmetric hopping regions. Within a critical phase, extended two-boson states can coexist with bounded boson pairs forming doublons, which could also subject to NHSEs under the open boundary condition (OBC). The borders between extended, critical and localized phases are further found to carry topologically nontrivial signatures.